# Signal Processing over Time-Varying Graphs: A Systematic Review

### v1

@@ IV-B Offline GSP Algorithms
1) General Offline GSP Methods: One naive way to process the time-varying signals is to apply the graph convolution in (4) or a GCN to the observations as we receive new observations of the graph signal. However, this method does not capture any temporal changes in the signal, which often results in poor performance. To effectively process time-varying signals on graphs, it is essential to capture both spatial and temporal dynamics.
@@
Another way to use GSP to process time-varying graph signals is to treat each time instance as a feature and formulate a multi-feature GSP cost function to minimize the error between estimation and ground truth :
@@
where is the Frobenius Norm, is the regularization weight, Tr is the trace, and is a spatial-smoothness-based regularization that minimizes the total variation of the results [5]. The solution can be found by calculating the gradient and setting it to zero. This method has 2 potential drawbacks. First, treating signals from different time instances as features means that the method is still inherently time-invariant, which does not take into account any temporal dynamics within the signal. Second, after some algebraic manipulations, the final solution requires the calculation of the matrix inverse, which can be unstable and unobtainable when the number of nodes in a graph is very large. This first limitation can potentially mitigated by replacing the spatial smoothness in (16) with the Sobolev Smoothness, which captures the smoothness in both the spatial domain and the temporal domain [37]. To address the second limitation, one can approximate the matrix inverse using Newton’s method, which gives us an iterative method that can be implemented in a distributed manner [85].
@@
2) Joint Time and Graph Fourier Transform: The foundational work on graph wavelets [6] laid the groundwork for localized spectral analysis on static graph signals, which later evolved into more comprehensive spatiotemporal frameworks. To handle scenarios where graph signals evolve over time, it is essential to consider both spatial and temporal dimensions simultaneously using GSP approaches. This has led to the development of Time-vertex methods and the Joint Time and Graph Fourier Transform (JFT) that integrate time-domain information directly into the spectral operations seen in the GSP framework [86]. The Joint Time-Vertex Fourier Transform captures both the spatial domain variations and the temporal domain variations by combining GFT and DFT, providing a comprehensive analysis of spatiotemporal signals, merging classical signal processing and GSP into a unified framework [87, 88]. The JFT also made the definition of dynamic wavelets on graphs possible for tracking the evolving time-varying graph signals [89]. This allows it to effectively decompose signals into joint time-vertex frequency components, offering insights into both temporal dynamics and graph-structured relationships. To ensure stability and enhance performance, an additional regularization term is introduced, promoting smoothness across both graph and time dimensions [88].
@@
3) Time-series Analysis on Graphs: As we explained earlier, conventional multivariate time series models, such as Vector Autoregressive (VAR) and Vector Autoregressive Moving Average (VARMA) processes, can be applied directly to graph signals since these signals are inherently multivariate [90]. However, conventional multivariate time series models are primarily designed for temporal tasks, which limits their ability to exploit the spatial relationships captured within the graph topologies. Taking the VARMA model as an example, it has the expression of
@@
where is the AR order, is the MA order, is a multi-variate random variable, is the by AR parameter, and is the by MA parameter, is an identity matrix, and denotes the matrix of white noise at time . It becomes difficult to accurately infer model parameters in (17) when there are nodes consistently missing, causing conventional VARMA models to underperform. To tackle temporal variability and represent space-time cross-dimensional interactions, one can merge time series analysis methods with GSP.
@@
Time-series analysis algorithms were introduced into GSP by specifying their parameters in the spectral domain via GFT, using the multivariate time-series as graph signals and potentially decreasing parameter size compared to traditional VAR and VARMA [91, 92]. The GSP analogy of VAR and VARMA is the graph Vector Autoregressive (GVAR) model [91] and the graph Vector Autoregressive–Moving-Average (GVARMA) model [93] respectively. Specifically, in VAR, the temporal association is modeled by an AR process, and the spatial association is captured by describing the matrix coefficients of the AR process as graph polynomial filters. In GVARMA, the VARMA coefficients are designed independently from the underlying graph but instead are as spectral filters. The filter design now takes into account concurrently both the spatial and temporal domains:
@@
Additionally, these parameters can also be defined using spatial graph methods by using the Chebyshev approximation (6) [92]:
@@
Parameter reduction in GVARMA can be achieved as follows. Even though both VARMA and GVARMA parameters have size by , the GVARMA in (19) defines the parameters in the spectral domain as shown in (19) by using the filter , which has only parameters. The number of parameters can be further reduced by defining a frequency set with , and then defining parameters only for the frequencies that present in . As a result, the actual size of the parameters that the GVARMA needs to learn in (19) is drastically reduced to . Similarly, if the GVARMA was to be derived from the spatial domain as shown in (20), the actual sizes of the parameters that the GVARMA needs to learn are and in and respectively, for all and .
@@
VARMA and GVARMA models are limited by their assumption of constant covariance, which makes them insufficient for capturing time-varying features such as volatility. Volatility, a key concept in finance, has been extensively studied by researchers. In time series analysis, models such as Autoregressive Conditional Heteroscedasticity (ARCH) [94] and its generalized version (GARCH) [95] were proposed to address symmetric volatility. To capture asymmetric volatility, the exponential GARCH (EGARCH) model [96] and the GJR-GARCH model [97] were subsequently developed. Recently, a unified framework for defining GARCH models on graphs was introduced [98]. The G-GARCH model shares the same model formula as GVARMA in (18) with the exception that the covariance matrix of also follows a GVARMA model instead of having constant variance. The G-GARCH model treats graph signals and their residuals as graph-stationary processes [99]. By leveraging the graph spectral domain, G-GARCH decomposes a multivariate GARCH model into a linear combination of several univariate GARCH processes, enabling efficient processing of graph-structured data. The G-GARCH model can also be derived into asymmetric models such as the G-EGARCH and G-GJR-GARCH [98].
V Graph Neural Networks for Time-Varying Graphs
Time-varying graphs can be effectively processed using signal-processing-based techniques, which offer powerful tools for analyzing and understanding graph-structured data. These GSP methods can be further enhanced by addressing certain challenges. They usually rely on priors for the parameters and filter design for simple and efficient implementation, but the priors are sometimes inaccessible. Additionally, augmenting traditional linear GSP techniques with mechanisms to capture complex non-linear patterns can significantly improve their ability to handle network data with intricate dependencies and dynamics.
With the recent revival of Neural Networks (NN) [100, 101, 102], Graph Convolutional Networks (GCN) was proposed, which combined nonlinear activation functions and learnable parameters with graph convolution operators in (6) and attained promising performance for many tasks over graphs, such as node classification, link prediction, and node regression [8, 10, 28, 103]. Despite the success of GCN in processing general static graph data, when the graph varies over time, the GCN cannot read the temporal information and fails to produce satisfying predictions. Therefore, developing graph neural networks that can adaptively learn and process time-varying graphs has been an increasingly attended issue. In this chapter, we summarize the existing works for TV-GSP and classify them into three technical directions: spatiotemporal neural network architectures, dynamic neural network architectures, and graph continual learning.
V-A Spatiotemporal Graph Neural Networks
@@
1) Decoupled Learning Methods: Learning spatial and temporal features respectively is the most natural idea for dealing with time-varying challenges. Based on this idea, several decoupled spatiotemporal graph neural networks are proposed [15, 104, 105]. The spatiotemporal Graph Convolutional Networks (STGCN) [15] is the most typical decoupled learning method, and it mainly comprises two parts for the sake of successive spatiotemporal learning, which can be expressed as:
where and separately represent the temporal and graph convolutional operators. They are defined as:
where and are two feature maps generated from by two 1-D convolution operators (i.e., one for feature representation and one for gating). can be the feature maps of the last neural network layer; represents the -th learnable parameter, and applies the -th order Chebyshev polynomial [10] to the graph Laplacian matrix.
@@
2) Integrated Learning Methods: Compared with successive learning spatial and temporal features by two independent modules, merging spatial and temporal feature learning into one layer can be more conducive to learning joint information. Motivated by this, a series of integrated spatiotemporal graph neural networks are proposed [106, 107, 108], where the GCN is usually employed as a spatial feature extractor in an RNN layer for joint spatiotemporal learning. The Diffusion Convolutional Recurrent Neural Networks (DC-RNN) proposed to insert a tailored diffusion convolution into each RNN layer [106]. The proposed diffusion convolution can be modeled as:
@@
where and are two independent parameters for the transition matrices of diffusion process and reverse one respectively. Then, the diffusion convolution with activation function is used to replace the matrix multiplications in the Gated Recurrent Units (GRU) [109] to construct a Diffusion Convolution GRU (DCGRU) as a basic unit in its sequential model. By this means, the temporal information and spatial information is processed collaboratively.
@@
3) Noise-Aware Methods: Some methods also take into account the noise-induced data missing in modeling, and the representative work is the spatiotemporal Graph Gated Recurrent Units (SG-GRU) [110] that are developed to make prediction temporally and complete the missing nodes simultaneously with the aid of GFT and GRU. To achieve this, it first calculates the graph interpolation operator with a given admissible graph sampling operator :
@@
where means the minimal singular value of the given matrix; is the sub-matrix of the eigenvector matrix with columns restricted to the indices associated with the frequencies indicated in ; is assumed as bandlimited.
By this constraint, the graph interpolation operator can be found and used to recover the corrupted graph signal after forecasting as:
@@
where and are predicted by GRU and Spectral GRU from respectively. The major drawback of this empirical approach is that the GFT only acts as a domain transformation and the performance relies only on the GRU, which is not adapted to the graphs and lacks interpretability. Besides, the hyper-parameters selection (e.g., , , and ) heavily depends on the prior knowledge of the graph.
@@
Remarks: Constructing a Spaito-Temporal Graph Neural Network (STGNN) to collaboratively learn and process the time-varying features is a direct way to cope with time-varying challenges. As shown in Fig. 5(a), the paradigm of STGNNs can be summarized as consisting of a pair of time modules (e.g., 1D convolution, RNN and GRU) and spatial modules (e.g., GCN, ChebNet, and GAT), and an extra spatiotemporal fusion module. The difference between various STGNNs generally appears in the architecture design of these modules and how spatial and temporal features are fused. Some methods also consider denoising and inpainting in the process of spatiotemporal graph learning [111].
V-B Dynamic Graph Neural Networks
@@
Most of the STGNNs are designed for node-level time-varying signal processing, while the temporal variation on higher-order signals (e.g., edges and topologies) are overlooked. Recently, Dynamic Graph Neural Networks (DGNN) were proposed to address this problem. In DGNNs, the time-varying graph is divided into two classes. One is the Discrete-Time Dynamic Graph (DTDG), and another is the Continuous-Time Dynamic Graph (CTDG). In DTDG, the dynamic graph is represented by a series of snapshots that can be the graph adjacency matrices and corresponding node feature vectors over time. In CTDG, the time-varying graph is represented by an initial graph and a sequence of graph evolution events (e.g., node insertion, edge addition, and edge deletion). By different dynamic graph modeling, various DGNNs are introduced. We then discuss them separately.
1) DTDG Methods: As illustrated by Fig. 5(b), the typical pipeline of DGNNs for processing DTDG is drawn upon the experience from the architecture of RNN and can be described as:
where represents the current state of timestamp , which can be learned from the current graph topology (i.e., the adjacency matrix); represents the hidden state at timestamp . and are two distinct neural networks. In general, is a GCN, and can be a linear projection layer. Compared to the STGNN that assumes fixed graph topology, in DTDG, the evolving graph topology is considered as variables to assist learning at every timestamp to accurately reflect the node relationship with time [112, 113].
Jointly learning the graph topology and node embedding spatiotemporally at each timestamp is sometimes awkward and not scalable when dealing with large-scale graphs. To relieve this issue, considering the advantage of the Auto-Encoder (AE) in terms of information compression [114], AE is also involved in processing DTDG. Particularly, DynGEM [115] generates the graph embedding sequentially and incrementally based on the encoded information from the last timestamp to enhance scalability and grant the model with inherent ability to capture temporal correlation between adjacent snapshots. On the other hand, dyngraph2vec [116] attempts to encode the graph snapshots within a short time window as a whole based on the temporal smoothness assumption to improve efficiency.

### v2

@@ 3.2.1 General Offline GSP
A naive way to process the time-varying signals is to apply the graph convolution in (4) or a GCN to the observations as we receive new observations of the graph signal. However, this method does not capture any temporal changes in the signal and often results in poor performance. It is essential to capture both spatial and temporal dynamics to effectively process time-varying signals on graphs.
@@
Another way to use GSP to process TVG signals is to treat each time instance as a feature and formulate a multi-feature GSP cost function to minimize the error between estimation and ground truth :
@@
where is the Frobenius Norm, is the regularization weight, Tr is the trace, and is a spatial-smoothness-based regularization that minimizes the total variation of the results [130]. The solution can be found by calculating the gradient and setting it to zero. This method has 2 potential drawbacks. First, treating signals from different time instances as features means that the method is still inherently time-invariant, which does not take into account any temporal dynamics within the signal. Second, after some algebraic manipulations, the final solution requires the calculation of the matrix inverse, which can be unstable and unobtainable when the number of nodes in a graph is very large. This first limitation can potentially be mitigated by replacing the spatial smoothness in (36) with the Sobolev Smoothness, which captures the smoothness in both the spatial domain and the temporal domain [49]. To address the second limitation, one can approximate the matrix inverse using Newton’s method, which gives us an iterative method that can be implemented in a distributed manner [220].
3.2.2 Joint Time and Graph Fourier Transform
@@
The foundational work on graph wavelets [60] laid the groundwork for spectral analysis on static graph signals, which later evolved into more comprehensive spatiotemporal frameworks. To handle scenarios where graph signals evolve, it is essential to consider both spatial and temporal dimensions simultaneously. This has led to the development of Time-vertex methods and the Joint Time and Graph Fourier Transform (JFT) that integrates time-domain information directly into the spectral operations seen in the GSP framework [110]. The Joint Time-Vertex Fourier Transform captures both spatial domain variations and temporal domain variations by combining GFT and DFT, providing a comprehensive analysis of spatiotemporal signals and merging classical signal processing and GSP [52, 85]. The JFT also made the definition of dynamic wavelets on graphs possible for tracking the evolving TVG signals [53]. This allows it to effectively decompose signals into joint time-vertex frequency components, offering insights into both temporal dynamics and graph-structured relationships. An additional regularization term is introduced to ensure stability and performance, promoting smoothness across both graph and time dimensions [85].
3.2.3 Time-series Analysis on Graphs
@@
As we explained earlier, conventional multivariate time series models, such as Vector Autoregressive (VAR) and Vector Autoregressive Moving Average (VARMA) processes, can be applied directly to graph signals since these signals are inherently multivariate [13]. However, conventional multivariate time series models are primarily designed for temporal tasks, which limits their ability to exploit the spatial relationships captured within the graph topologies. Taking the VARMA model as an example, it has the expression of
@@
where is the AR order, is the MA order, is a multi-variate random variable, is the by AR parameter, and is the by MA parameter, is an identity matrix, and denotes the matrix of white noise at time . It becomes difficult to accurately infer model parameters in (37) when there are nodes consistently missing, causing conventional VARMA models to underperform. To address temporal variability and capture space-time interactions, time series analysis is integrated with GSP.
@@
Time-series analysis algorithms were introduced into GSP by specifying their parameters in the spectral domain via GFT, using the multivariate time-series as graph signals and potentially decreasing parameter size compared to traditional VAR and VARMA [119, 76]. The GSP analogy of VAR and VARMA is the graph Vector Autoregressive (GVAR) model [119] and the graph Vector Autoregressive–Moving-Average (GVARMA) model [73], respectively. Specifically, in VAR, the temporal association is modeled by an AR process, and the spatial association is captured by describing the matrix coefficients of the AR process as graph polynomial filters. In GVARMA, the VARMA coefficients are designed independently from the underlying graph, but instead are spectral filters. The filter design now takes into account both the spatial and temporal domains:
@@
Additionally, these parameters can also be defined by spatial graph methods using the Chebyshev approximation (7) [76]:
@@
Parameter reduction in GVARMA can be achieved as follows. Even though VARMA and GVARMA parameters have size by , GVARMA (39) defines the parameters in the spectral domain as shown in (39) by using the filter , which has only parameters. The number of parameters can be further reduced by defining a frequency set with , and then defining parameters only for the frequencies that are present in . As a result, the actual size of the parameters the GVARMA needs in (39) is drastically reduced to . Similarly, if the GVARMA was to be derived from the spatial domain as shown in (40), the actual sizes of the parameters that the GVARMA needs to learn are and in and respectively, for all and .
@@
VARMA and GVARMA models are limited by their assumption of constant covariance, which makes them insufficient for capturing time-varying features such as volatility. Volatility, a key concept in finance, has been extensively studied by researchers. In time series analysis, models such as Autoregressive Conditional Heteroscedasticity (ARCH) [40] and its generalized version (GARCH) [12] were proposed to address symmetric volatility. To capture asymmetric volatility, the exponential GARCH (EGARCH) model [128] and the GJR-GARCH model [51] were subsequently developed. Recently, a unified framework for defining GARCH models on graphs was introduced [63]. The G-GARCH model shares the same model formula as GVARMA in (38), with the exception that the covariance matrix follows a GVARMA model instead of constant variance. By leveraging the graph spectral domain, G-GARCH decomposes a multivariate GARCH model into a linear combination of several univariate GARCH processes, enabling efficient processing of graph-structured data. The G-GARCH model can also be derived into asymmetric models such as the G-EGARCH and G-GJR-GARCH [63].
Remarks Offline algorithms are most preferred when the long temporal sequences in dataset is available prior to processing and when long-term temporal dependencies need to be exploited. By accessing the longer temporal horizon, offline TVGSP methods can leverage richer information compared with their online counterparts. However, this capability comes with increased computational and memory requirements, and offline processing may be unsuitable in scenarios that demand immediate decision making or involve rapidly evolving graph signals. Therefore, offline TVGSP algorithms are best suited for applications prioritizing accuracy, global temporal understanding, and algorithmic expressiveness over real-time responsiveness.
4 Neural Networks for Time-Varying Graphs
Augmenting GSP models with nonlinear and adaptive components, instead of redesigning them from scratch, offers a principled path to handling richer dependencies and temporal dynamics. Early GNNs follow this philosophy by combining graph polynomial filters such as those in (7) with trainable weights and nonlinear activation functions, achieving strong performance on tasks including node classification, link prediction, and node regression [16, 34, 86, 21]. However, these GNNs are designed for static graphs. When applied to time-varying signals, they usually operate on a single static snapshot, largely ignoring the temporal dependencies in the data. As a result, their performance can degrade on tasks that rely on temporal structure, such as forecasting or dynamic pattern detection. This limitation has motivated a rich body of work on GNN architectures that explicitly model and adapt to TVG signals and topologies. It is worth noting that this section does not aim to provide an exhaustive survey of dynamic GNNs; instead, we highlight representative models that are most relevant to time-varying GSP and the perspective developed in this work. For completeness, we also discuss continual-learning-based DGNNs and CTDG methods. We emphasize that corresponding developments in TVGSP remain scarce; our goal here is therefore not to establish a one-to-one correspondence, but to identify important problem regimes where TVGSP may be extended in the future.
4.1 Spatiotemporal Graph Neural Networks (STGNNs)
4.1.1 Decoupled Learning Methods
@@
Learning spatial and temporal features, respectively, is the most natural idea when dealing with TVGs. Based on this idea, several decoupled spatiotemporal GNNs are proposed [207, 216, 217, 150]. The Spatiotemporal Graph Convolutional Network (STGCN) [207] is a typical decoupled learning method, in which spatial and temporal features are processed successively. The layer-wise propagation rule can be represented as
where and denote temporal and graph convolution operators, respectively, is the non-linear activation function, and is the spatiotemporal feature tensor at layer . The input to the first layer is typically a temporal window of graph signals, , where collects the node signals at time for an -node graph, and denotes a historical time interval used for training.
Each temporal operator is implemented as a gated temporal convolution. Specifically, is rooted from two parallel 1D (non-graph) temporal convolutions that produce a feature map and a gating map . The final temporal output is then obtained via element-wise modulation . For the graph component, STGCN adopts the spectral graph convolution framework, where the learnable filter is parameterized as a polynomial of the normalized Laplacian. In particular, is implemented using a Chebyshev polynomial approximation exactly as shown in (7), yielding an efficient and localized -hop filtering operation. Combining these instantiations gives
where and denote the two temporal feature maps, is the normalized Laplacian, are learnable coefficients, and denotes the -th order Chebyshev polynomial [34]. In STGCN, the temporal operators and in (41) take this gated-convolution form commonly found in non-graph NNs, while corresponds to the Chebyshev graph filter. Unlike the adaptive TVGSP filters discussed in Section 3, which can operate online as new observations arrive, STGCN and related STGNNs are typically trained offline on historical windows and then used for prediction under an implicit (quasi-)stationarity assumption, highlighting a key conceptual difference between learning-based STGNNs and adaptive GSP-based approaches.
From a GSP viewpoint, the spatial convolution in (42) is equivalent to a graph spectral filtering operator,
which directly parallels the Chebyshev polynomial approximation in (7). This interpretation establishes a direct theoretical correspondence between STGNN and GSP, showing that the spatial component of STGCN can be explicitly interpreted as a learnable GSP filter in the spectral domain, thereby providing interpretability while retaining neural flexibility.
4.1.2 Integrated Learning Methods
@@
Compared with the successive learning of spatial and temporal features by two independent modules, merging spatial and temporal feature learning into one layer can be more conducive to learning joint information. Motivated by this, a series of integrated spatiotemporal GNNs are proposed [103, 187, 72], where a GCN interpretable by GSP in (7) is employed as a spatial feature extractor in an RNN layer for joint spatiotemporal learning. For example, the Diffusion Convolutional Recurrent Neural Networks (DC-RNN) was proposed to insert a tailored diffusion convolution into each RNN layer [103]:
@@
where and are two independent parameters for the transition matrices of diffusion process and reverse one respectively. Then, the diffusion convolution with activation function is used to replace the matrix multiplications in the Gated Recurrent Units (GRU) to construct a Diffusion Convolution GRU (DCGRU) as a basic unit in its sequential model, processing temporal information and spatial information collaboratively. From a GSP viewpoint, this corresponds to 2 separate graph filtering operations representing the forward and reverse diffusion processes
which makes the spatial processing in DC-RNN explicitly interpretable as a learnable spectral filtering process, while the recurrent structure simultaneously models temporal evolution. This unified interpretation highlights how integrated STGNNs inherently perform joint GSP filtering and temporal state modeling.
4.1.3 Noise-Aware Methods
@@
Some methods also take into account the noise-induced data missing in modeling, and the representative work is the spatiotemporal Graph Gated Recurrent Units (SG-GRU) [96] that are developed to make predictions temporally and complete the missing nodes simultaneously with the aid of GFT and GRU. To achieve this, it first calculates the graph interpolation operator with a given admissible graph sampling operator :
@@
where means the minimal singular value of the given matrix; is the sub-matrix of the eigenvector matrix with columns restricted to the indices associated with the frequencies indicated in ; is assumed as bandlimited. By this constraint, the graph interpolation operator is used to recover the corrupted graph signal after forecasting:
@@
where MLP stands for multi-layer perceptron, and are predicted by GRU and Spectral GRU from respectively. The major drawback of this empirical approach is that the GFT only acts as a domain transformation, and the performance relies only on the GRU, which is not adapted to the graphs and lacks interpretability. Besides, the hyperparameters selection (e.g., , , and ) heavily depends on the prior knowledge of the graph.
@@
Remarks: Constructing an STGNN to collaboratively learn and process the time-varying features is a direct way to cope with time-varying challenges. The STGNNs mostly resemble the TVGSP approaches, such as the spatial graph convolution seen in (8) or (5), with the inclusion of learnable parameters and activation functions. As shown in (a) of Fig. 2, the paradigm of STGNNs can be summarized as consisting of a pair of time modules (e.g., 1D convolution, RNN and GRU) and spatial modules (e.g., GCN, ChebNet, and GAT), and an extra spatiotemporal fusion module. From a GSP perspective, the spatial modules in these GNNs can be interpreted as graph filtering operations implemented via the GFT, where learned parameters implicitly define spectral responses over the graph. The difference between various STGNNs generally appears in the architecture design of these modules and how spatial and temporal features are fused. Differences across STGNN variants primarily arise from how these modules are architected and how spatial and temporal information is fused, while some approaches additionally incorporate denoising or inpainting mechanisms during spatiotemporal learning [20]. This close relationship suggests promising opportunities for transferring GSP insights on temporal regularization, filter design, and robustness into the design and interpretation of dynamic GNNs.
4.2 Dynamic Graph Neural Networks
@@
Most of the STGNNs are designed for node-level time-varying signal processing, while the temporal variation on higher-order signals (e.g., edges and topologies) is overlooked. Recently, Dynamic Graph Neural Networks (DGNN) were proposed to address this problem. In DGNNs, the TVG is divided into DTDGs and CTDGs, the two classes of dynamic graphs that correspond to the categories we defined in Section 2. By different dynamic graph modeling, various DGNNs are introduced. We then discuss them separately.
4.2.1 DTDG Methods
As illustrated in Fig. 2(b), typical DGNNs for DTDG processing follow an RNN-inspired pipeline. At each discrete timestamp , a spatial GNN extracts node representations from the current graph snapshot , and a temporal module propagates information over time, i.e.,
where denotes the input graph signal (node features), denotes the learned node representations at time , and is a hidden state capturing temporal dependencies. In practice, is typically a GCN-inspired structure implemented using , while is often implemented as a recurrent or linear projection module. Unlike STGNNs with fixed topology, DTDG models explicitly leverage the evolving graph structure at each timestamp to better capture time-varying node relationships [115, 97]. EvolveGCN is another DTDG-GNN framework in which an RNN dynamically updates the GCN parameters across discrete graph snapshots, enabling effective learning on evolving DTDGs [133]. Variational Graph Recurrent Neural Networks model TVGs as DTDGs using a variational autoencoder framework, in which latent node representations evolve through a recurrent prior across successive graph snapshots, enabling effective representation learning and link prediction [58]. These models are representative of GNNS using DTDGs; here, we highlight them as key examples, while recognizing that the broader DGNN literature is considerably richer.
From a GSP perspective, the function in the above DTDG pipeline can be interpreted as applying a TVG filter to the node features at each time step. In particular, our GSP formulation in Section 3 considers time-varying spatial operators of the form (28). When such an operator is used as the spatial component of , a natural GSP-inspired realization is
where denotes the node features at time and is a pointwise nonlinearity. In this view, many DTDG GNNs implicitly learn a sequence of TVG filters through their parameters, whereas the GSP-based formulation in (28) makes these filters explicit in terms of powers of the time-varying Laplacian that are directly determined by the evolving topology . This correspondence provides a principled bridge between DTDG GNN architectures and TVGSP, and suggests opportunities to transfer stability and regularization insights from GSP to the design of dynamic GNNs.
4.2.2 CTDG Methods